Circular No. 7091

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999E IN ANONYMOUS GALAXY E. Cappellaro and M. Turatto, Padova Observatory; and P. Mazzali,

Trieste Observatory, write: “Based on the spectrum taken by Jha et al.,

given at http://cfa-www.harvard.edu/cfa/oir/Research/supernova/

spectra/sn99e.99Jan19.gif, and on the description on IAUC 7090 of the Filippenko spectrum, SN 1999E is very similar to SN 1997cy (Turatto et al., in preparation). Although the evolution is very slow, the best match is with the spectrum taken ∼ 4 months after discovery (http://athena.pd.astro

.it/~supern/news/sn97cy 251198.gif), with emission lines of H, Ca II,

and [Fe II] resulting from the interaction of the ejecta with a dense circumstellar material, as in SN 1988Z (Turatto et al. 1993, MNRAS 262, 128). SN 1997cy has been associated with GRB 970514 (Woosley et al.,

Ap.J., submitted, and http://xxx.lanl.gov/abs/astro-ph/?9806299),

it reached M_{v} > = −20.06 (H_{0} = 65 km s^{−1} Mpc^{−1}), and it was to our knowledge the intrinsically brightest supernova ever discovered. Based on the analogy with 1997cy, we expect that the spectral evolution and the luminosity decline of 1999E will both be very slow.”

A. V. Filippenko, D. C. Leonard, and A. G. Riess, University of California at Berkeley, report: “Another spectrum (range 380–630 nm) of the luminous, peculiar type-IIn supernova 1999E (cf. IAUC 7089, 7090) was obtained on Jan. 20 UT with the Keck-II telescope. The features are similar to those reported by Jha et al. Combining both of our own spectra, perhaps the closest match is to that of SN 1997cy (whose spectrum was obtained by B. P. Schmidt), which was likewise extremely luminous. In some respects (e.g., breadth), the undulations also resemble those of the luminous SN 1998bw, which may have been associated with GRB 980425.”

SUPERNOVAE 1998fc AND 1998fd IN ANONYMOUS GALAXIES Filippenko et al. also report that a spectrum (range 510–880 nm) ob-

tained on Jan. 20 UT with the Keck-II telescope reveals that SN 1998fc (cf. IAUC 7082) is of type Ia at redshift 0.10, perhaps a month past maximum brightness. The emission-line intensity ratios are somewhat peculiar; specifically, the feature at rest wavelength 550 nm is unusually strong. A similar spectrum obtained on Jan. 19 reveals that SN 1998fd (cf. IAUC 7082) is of type Ia, a few weeks past maximum brightness; a redshift of 0.24 was measured from absorption lines in the host galaxy.

1999 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green